# Pipeline Gradient-based Model Training on Analog In-memory Accelerators

### v1

@@ Abstract
Aiming to accelerate the training of large deep neural models in an energy-efficient way, an analog in-memory computing (AIMC) accelerator emerges as a solution with immense potential. In AIMC accelerators, trainable weights are kept in memory without the need to move from memory to processors during the training, significantly reducing overhead. However, although the in-memory feature enables efficient computation, it also constrains the use of data parallelism since copying weights from one AIMC to another is expensive. To enable parallel training using AIMC, we propose synchronous and asynchronous pipeline parallelism for AIMC accelerators inspired by the pipeline in digital domains. This paper provides a theoretical convergence guarantee for both synchronous and asynchronous pipelines in terms of both sampling and clock cycle complexity, which is non-trivial since the physical characteristic of AIMC accelerators leads to analog updates that suffer from asymmetric bias. The simulations of training deep models on real datasets verify the efficiency of pipeline analog training.
1 Introduction
Deep Neural Networks (DNNs) have facilitated remarkable progress in a wide range of applications. As the model sizes and dataset volumes grow exponentially, DNN training and inference become time-consuming and resource-intensive. For example, it requires 184 thousand GPU hours to train an LLAMA2 7 billion model, and this time increases to 1.7 million GPU hours for its 70 billion version [1]. Stemming from the escalating computational demands of large-scale DNN training and inference, the pursuit of specialized hardware accelerators emerges. In principle, ideal specialized hardware for DNN should be efficient, support large-scale parallelism, and remain lower energy-economic simultaneously. As the demand for efficient processing intensifies, the fundamental backbone of both training and inference largely relies on matrix-vector multiplication (MVM) operation, which is the core component of the ubiquitous fully connected and convolutional layers. Typically, each forward and backward pass requires on the order of millions of multiply-accumulate (MAC) operations, necessitating specialized hardware with high performance and low energy consumption.
The promise of analog computing. To this end, a promising solution to accelerate the MVM operations is the analog in-memory computating (AIMC) accelerator, wherein both the input and output of MVM operation are analog signals, and the trainable weights of DNN layers are stored in analog crossbar tiles [2]. Different from traditional Von Neumann accelerators like GPU and TPU, which require weight movement from memory to processor before computation, the AIMC accelerator keeps the weight stationary and performs MVM operations in memory. Without moving weights frequently, AIMC accelerators enable time- and energy-efficient MVM computations; see Figure 1.
The curse of data parallelism. While efficient computation is a benefit, in-memory computing imposes limitations on scalability. Specifically, contemporary DNN optimizers depend on mini-batch SGD to facilitate highly parallel training. In the Von Neumann computation architecture, the most efficient parallel mechanism is data parallelism [3], where DNN weights are replicated to different computation devices to support gradient descent with large batches [4, 5]. On the contrary, in AIMC architecture, the trainable weights are stored in the analog tiles in the form of device conductances, and thus, copying weights from one analog tile to another is expensive and prone to error. Therefore, data parallelism, the enabling factor of large-scale AI training, is unrealistic for AIMC accelerators.
Due to the physical limitations of the AIMC accelerator, we instead resort to pipeline parallelism [6], another prevalent parallel mechanism. In pipeline parallelism, a large model is partitioned into a series of smaller stages and gets distributed across multiple devices, with each stage corresponding to the computation of a few layers. After a device finishes the process of one datum (or one batch), it passes the intermediate outputs to the next devices and begins the process of the next data. In this way, the model that is too large to fit into one device can be separated into different AIMC devices, enlarging the efficient memory. Meanwhile, the computation on different devices can happen simultaneously, and the systems can process more data in a given period of time.
While pipelining is a simple idea, implementing it in analog training faces a series of crucial challenges. First, AIMC accelerators usually suffer from noisy gradient signals and asymmetric updates [7, 8], which introduce errors on all forward, backward, and update passes. Second, to fully realize the potential of pipeline parallelism, the weight update pass usually involves stale gradients. Both of them incur errors in the training, leading to the following question:
How to design analog pipeline parallel training approaches that do not impair the training?
This paper answers this question by analyzing the sample complexity of different pipeline algorithms and wall-clock time. Before that, we will briefly review prior work that attempted to solve the issues.
@@ 1 Introduction
Figure 1: (Left) Illustration of MVM computation in AIMC accelerators. The weight $W^{(m)}$ on layer $m$ is stored in a crossbar tile consisting of an array of resistors. The $(i,j)$-th element of $W^{(m)}$ is represented by the conductance of the $(i,j)$-th resistor. To perform an MVM operation $z^{(m)}=W^{(m)}x^{(m)}$, voltage $[x^{(m)}]_{j}$ is applied between $j$-th and $(j+1)$-th row. By Ohm’s law, the current is $I_{ij}=[W^{(m)}]_{ij}[x^{(m)}]_{j}$; and by Kirchhoff’s law, the total current on the $i$-th column is $\sum_{j}I_{ij}=\sum_{j}[W^{(m)}]_{ij}[x^{(m)}]_{j}$. Unlike the digital counterpart, no movement of $W^{(m)}$ is required. (Right) Speedup of the proposed pipeline training methods as the number of stages increases compared with single machine training case (stage=1); see Section 5 for more details.
Gradient-based training on AIMC accelerators. Analog SGD implemented by pulse update is extremely efficient [8]. However, the asymmetric update and noisy gradient lead to large asymptotic errors in training [7]. To overcome these issues, a family of analog training algorithms, Tiki-Taka (TT) has been proposed. TT-v1 [9] heuristically introduces an auxiliary analog array as a gradient approximation to eliminate the asymptotic error. To demonstrate its efficiency, [10] establishes a model to characterize the dynamic of the analog training and provides a theoretical convergence guarantee. To filter out the high-frequency noise introduced by inter-device transfer, TT-v2 [11] introduces a digital array. Based on TT-v1/v2, Chopped-TTv2 (c-TTv2) and Analog Gradient Accumulation with Dynamic reference (AGAD) [12] are proposed to mitigate the impact of inaccurate symmetric point correction [13]. Despite the above efforts, analog training of large models on huge datasets is still a major challenge.
Synchronous and asynchronous parallelism. Typically, existing pipeline designs can be classified into two categories. If the weights are updated using the latest gradients, it is synchronous pipeline; otherwise, it is asynchronous pipeline. As the computing requirement of training large models emerges, a plethora of pipeline parallel libraries has been developed, including Gpipe [14, 15], Deepspeed [16], FSDP [17], Megetron-LM [18], Colossal-AI [19], just to name a few. Nevertheless, they are synchronous pipeline that enjoys the same dynamic as SGD, so they have the same convergence guarantee as SGD. Apart from that, researchers also made some effort to explore the asynchronous pipeline. For instance, [20] studied the asynchronous pipeline setting and proposed weight-prediction approaches to compensate for the delay. PipeDream [21] combined the synchronous and asynchronous pipeline parallelisms to achieve greater speedup. To mitigate the negative effect of stale gradients, [22] and [23] have proposed only overlapping gradient computation and communication so that delay is always one. However, these algorithms are not for analog pipeline training. Researchers also attempted to design analog pipeline circuits to accelerate the training [24, 25], but they focus on the synchronous pipeline. On the theory side, there are a few studies on pipeline training on digital devices such as [26], but these results rely on weight stashing, which is expensive for analog accelerators. As [21] points out, delay affects both forward and backward passes in the asynchronous pipeline training. It renders the increment no longer a valid gradient without the weight stashing and makes the analysis challenging.
In the context of prior work, we summarize our key contributions as follows.
@@ 1. C1)
    To fully embrace the promise of analog computing, we systematically study the pipeline training on AIMC accelerators, which enables training larger models by distributing models across multiple accelerators and increasing computation density. We study the existing work and classify them as synchronous pipelines, which achieve considerable speedup, but their computation density is still sub-optimal. To further increase the computation density, we propose the first asynchronous pipeline training algorithm on AIMC accelerators.
@@ 2. C2)
    We provide a convergence guarantee for the proposed algorithms. Theoretical results show that two strategies have pros and cons: synchronous pipeline has a better complexity while its computation density in the device is lower; asynchronous pipeline reduces device idleness but involves stale gradients. The main results are summarized in Table 1.
@@ 3. C3)
    We conduct empirical simulation to demonstrate the efficiency of pipeline training. The results show that both pipeline algorithms significantly outperform their non-pipeline counterparts. Despite a slightly larger sample complexity, the asynchronous pipeline achieves the same accuracy with fewer clock cycles because of its high computation density. Furthermore, the asynchronous pipeline achieves linear-speedup at least within the range of 1-8 devices; see Figure 1.
Pipeline Type Sample Comp. Comp. Density Clock Cycle
w.o. Pipeline $O\left(\sigma^{2}\varepsilon^{-2}\right)$ $\frac{1}{M}$ $O\left(M\sigma^{2}\varepsilon^{-2}\right)$
Synchronous [Theorem 1] $O\left(\sigma^{2}\varepsilon^{-2}\right)$ $\frac{B}{M+B-1}$ $O\left((M+B)B^{-1}\sigma^{2}\varepsilon^{-2}\right)$
Asynchronous [Theorem 2] $O\left(\sigma^{2}\varepsilon^{-2}+\varepsilon^{-1}\right)$ $1$ $O\left(\sigma^{2}\varepsilon^{-2}+\varepsilon^{-1}\right)$
Table 1: Comparision of different pipeline strategies. Sample complexity (Sample Comp.) is the number of samples needed for achieving training loss $\leq\varepsilon$. Computation density (Comp. Density) is the time that is the portion of time when the system is computing the gradients. In the table, $M$ is the pipeline depth, and $B$ is the micro-batch size. Clock Cycle measures the running time for achieving training loss $\leq\varepsilon$, which in math is sample complexity divided by clock cycle.
2 Model Training on AIMC Accelerators
In this section, we first introduce the generic gradient-based training algorithm, encompassing digital algorithms running on devices like GPU and TPU, and algorithms running on AIMC accelerators. Consider a standard model training problem on a data distribution ${\mathcal{D}}$, given by
$W^{*}:=\operatornamewithlimits{arg\,min}_{W\in{\mathbb{R}}^{D}}f(W)~{}~{}~{}{\rm with}~{}~{}~{}f(W):={\mathbb{E}}_{\xi\sim{\mathcal{D}}}[f(W;\xi)]$ (1)
@@ 2 Model Training on AIMC Accelerators
where $W$ is all trainable weights of the model vectorized into a $D$ dimensional vector.
@@
DNN model. Consider the multi-layer DNN, which consists only of linear layers and activation functions. Let $[M]:=\{1,2,\cdots,M\}$ denote a positive integer set. The trainable weights are separated into $M$ different blocks, i.e., $W=[W^{(1)},W^{(2)},\cdots,W^{(M)}]$, where $W^{(m)},\forall m\in[M]$ is a matrix. The training of DNN typically consists of three phases: forward, backward, and update. Suppose $x$ and $y$ are the feature and label of data $\xi=(x,y)\in{\mathcal{X}}\times{\mathcal{Y}}$, where ${\mathcal{X}}$ and ${\mathcal{Y}}$ are feature and label space, respectively. The forward pass of the model for layer $m\in[M]$ can be expressed by
$\textsf{Forward}\qquad x^{(1)}=x,~{}~{}~{}~{}$ $z^{(m)}=W^{(m)}x^{(m)},~{}~{}~{}~{}x^{(m+1)}=g^{(m)}(z^{(m)})$ (2)
@@
where $g^{(m)}$ is the activation function of layer $m$, and $x^{(M)}$ is the feature from the last layer, which is fed together with label $y$ into the loss function $\ell(\cdot,\cdot):{\mathcal{Y}}\times{\mathcal{Y}}\to{\mathbb{R}}$ and yields $f(W;\xi)=\ell(x^{(M+1)},y)$. For the backward pass, define the backward error $\delta^{(m)}:=\frac{\partial f(W;\xi)}{\partial z^{(m)}}$, wherein the error of last layer is first computed by $\delta^{(M)}=\nabla_{x^{(M+1)}}\ell(x^{(M+1)},y)\nabla g^{(M)}(z^{(M)})$. Subsequently, each layer $m\in[M-1]$ computes the next recursion according to
@@
In the forward pass, the activation $z^{(m)}$ is stashed until the backward signal $\delta^{(m+1)}$ returns. Typically, the computation $\nabla g^{(m)}(z)$ is light-weight, and hence the major overhead of the backward pass lies at the MVM among $W$ and $\delta$. With the forward and backward signals, the gradient with respect to the weight $W^{(m)}$ of layer $m$ is computed by
@@
where $\otimes$ is the outer product between two vectors. Having the gradients of each layer, we can use different gradient-based algorithms to update the weights.
Gradient-based training on AIMC accelerators. After getting the backward signal $\delta^{(m)}$, gradient-based algorithms can be used to update the weights. In digital devices, one of the most popular training algorithms is SGD (SGD), wherein the matrix $\nabla_{W^{(m)}}f(W;\xi)=\delta^{(m)}\otimes x^{(m)}$ is computed and added to $W^{(m)}_{k}$, written by $W_{k+1}^{(m)}=W_{k}^{(m)}-\alpha\nabla_{W^{(m)}}f(W_{k};\xi_{k})$; $\alpha>0$ is a positive learning rate. In the paper, the subscript $k$ is used to indicate the index of gradient update.
Different from digital devices, AIMC accelerator uses rank-update to update the gradient update, which constructs a series of pulses and sends them to its analog array [8]. Without computing the whole matrix $\delta^{(m)}\otimes x^{(m)}$ explicitly, AIMC accelerators perform the update operation in a fast and energy-efficient way. Nevertheless, rank-update relies on the special structure that the gradient of $W^{(m)}_{k}$ is a rank-one matrix, which brings us further challenges.
On top of that, as existing work pointed out [9, 10], the update on AIMC accelerators suffers from asymmetric bias, and thus the dynamic of Analog SGD in single-machine training can be written as
$W_{k+1}^{(m)}=$ $\ W_{k}^{(m)}-\alpha\nabla_{W^{(m)}}f(W_{k};\xi_{k})-\frac{\alpha}{\tau}|\nabla_{W^{(m)}}f(W_{k};\xi_{k})|\odot W_{k}^{(m)}$ (5)
where $\tau$ is a device-dependent constant. When $\tau\rightarrow\infty$, Analog SGD reduces to Digital SGD.
The main challenge coming from the asymmetric is the saturation phenomenon [9]. Consider the update of $(i,j)$-th element of matrix $W_{k}^{(m)}$, whose update is written as
$([W_{k+1}^{(m)}]_{ij}-[W_{k}^{(m)}]_{ij})^{2}=$ $\ \alpha^{2}(\text{sign}([\nabla_{W^{(m)}}f(W_{k};\xi_{k})]_{ij})-[W^{(m)}_{k}]_{ij}/\tau)^{2}(\nabla_{W^{(m)}}f(W_{k};\xi_{k}))^{2}$
$\approx$ 0if [$\ 0~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\text{if }[W^{(m)}_{k}]_{ij}\approx\tau~{}\text{sign}([\nabla_{W^{(m)}}f(W_{k};\xi_{k})]_{ij})$ (6)
which means that the increment is suppressed when $[W^{(m)}_{k}]_{ij}$ gets close to the $\tau$. Therefore, we usually introduce a proper scaling coefficient so that the saturation never happens during the training.
@@
Figure 2: Illustration of pipelines with 4 devices ($M=4$). Each mini-batch is split into $B=4$ micro-batches. Each color represents one micro-batch, and each row from bottom to top represents stages 0 to 4. Each column corresponds to a clock cycle in which one micro-batch is processed. The white square indicates the idle device. (Top) Vanilla model parallelism without pipeline. All weights are updated using gradient-based algorithms after the computation of one mini-batch is completed, which is not presented in the figure. (Middle) Synchronous pipeline with micro-batch count $B=5$. The update happens after all 5 micro-batches are processed. (Bottom) Asynchronous pipeline. The update happens once the gradient of one micro-batch is achieved. The grey square indicates the corresponding device is processing data that is not fully reflected in the figure.
3 Analog Pipeline Model Training
This section introduces pipeline parallelism and shows how it increases computational efficiency.
Consider training a large model using a mini-batch size $B_{\text{mini}}$, which is too large to fit into one device and motivates us to adopt model parallelism. In model parallelism, the computation is partitioned into a sequence of stages, which, in general, consists of a consecutive set of layers in the model. Each stage is mapped on one digital or analog device. For simplicity, we assume each stage contains one layer only. Apart from the model separation, a mini-batch of data is also split into $B$ micro-batch of size $B_{\text{micro}}:=B_{\text{mini}}/B$. For simplicity, let the micro-batch size $B_{\text{micro}}=1$ in Sections 3-4 so that the subscript $b$ in $\xi_{k,b}$ can represent the index of micro-batch or data, and $B=B_{\text{mini}}$ can represent the number of micro-batch and mini-batch size $B_{\text{mini}}$. It could be easily extended to general cases.
Clock cycle. We define the maximum time for each stage to process one micro-batch as the clock cycle. Ideally, all devices would process a micro-batch in the same amount of time, avoiding idle periods during a clock cycle. Although achieving this is challenging, various task scheduling approaches have been proposed to balance computation [21, 14]. To explore the potential of pipeline training, this paper uses the clock cycle and stage as the minimal time and basic scheduling units, respectively. Based on scheduling methods, there are three types of model parallelism: vanilla model parallelism without pipelining, synchronous pipelining, and asynchronous pipelining.

### v2

# On the Convergence Theory of Pipeline Gradient-based Analog In-memory Training

@@ Abstract
Aiming to accelerate the training of large deep neural networks (DNN) in an energy-efficient way, analog in-memory computing (AIMC) emerges as a solution with immense potential. AIMC accelerator keeps model weights in memory without moving them from memory to processors during training, reducing overhead dramatically. Despite its efficiency, scaling up AIMC systems presents significant challenges. Since weight copying is expensive and inaccurate, data parallelism is less efficient on AIMC accelerators. It necessitates the exploration of pipeline parallelism, particularly asynchronous pipeline parallelism, which utilizes all available accelerators during the training process. This paper examines the convergence theory of stochastic gradient descent on AIMC hardware with an asynchronous pipeline (Analog-SGD-AP). Although there is empirical exploration of AIMC accelerators, the theoretical understanding of how analog hardware imperfections in weight updates affect the training of multi-layer DNN models remains underexplored. Furthermore, the asynchronous pipeline parallelism results in stale weights issues, which render the update signals no longer valid gradients. To close the gap, this paper investigates the convergence properties of Analog-SGD-AP on multi-layer DNN training. We show that the Analog-SGD-AP converges with iteration complexity ${\mathcal{O}}\left(\varepsilon^{-2}+\varepsilon^{-1}\right)$ despite the aforementioned issues, which matches the complexities of digital SGD and Analog SGD with synchronous pipeline, except the non-dominant term ${\mathcal{O}}(\varepsilon^{-1})$. It implies that AIMC training benefits from asynchronous pipelining almost for free compared with the synchronous pipeline by overlapping computation.
I Introduction
@@
Figure 1: Illustration of MVM computation in AIMC accelerators. The weight $W^{(m)}$ on layer $m$ is stored in a crossbar tile. The $(i,j)$-th element of $W^{(m)}$ is represented by the conductance of the $(i,j)$-th resistive element. MVM operation $z^{(m)}=W^{(m)}x^{(m)}$ is performed by applying voltage $x^{(m)}]_{j}$ between $j$-th and $(j+1)$-th row. By Ohm’s law, the current is $I_{ij}=[W^{(m)}]_{ij}[x^{(m)}]_{j}$; and by Kirchhoff’s law, the total current on the $i$-th column is $\sum_{j}I_{ij}=\sum_{j}[W^{(m)}]_{ij}[x^{(m)}]_{j}$. Unlike the digital counterpart, no movement of $W^{(m)}$ is required for MVM calculation.
Deep Neural Networks (DNNs) have facilitated remarkable progress in a wide range of applications. As the model sizes and dataset volumes grow exponentially, DNN training and inference become time-consuming and resource-intensive. For example, training an LLAMA2 7B model requires 184 thousand GPU hours, and this increases to 1.7 million GPU hours for its 70B version [1]. In response to the escalating computational demands of large-scale DNN training and inference, the pursuit of specialized hardware accelerators has emerged. The ubiquitous fully connected and convolutional layers, whose core operation is matrix-vector multiplication (MVM), form the fundamental backbone of both training and inference as the demand for efficient processing intensifies. Typically, each forward and backward pass requires millions of multiply-accumulate (MAC) operations, necessitating specialized hardware with high performance and low energy consumption.
The promise of analog computing. To this end, a promising solution to accelerate the MVM operations is the analog in-memory computing (AIMC) accelerator, wherein both the input and output of the MVM operation are analog signals like current or voltage. The DNN weights are stored in analog crossbar tiles consisting of resistive elements [2] where the conductance of the resistive element represents the weight magnitude. Unlike traditional Von Neumann accelerators such as GPUs and TPUs, which move weights from memory to the processor during computation, AIMC accelerators keep the weights stationary and perform MVM operations in memory. Without moving weights frequently, AIMC accelerators enable time- and energy-efficient MVM computations; see Fig. 1.
The curse of data parallelism. While efficient computation is beneficial, in-memory computing imposes scalability limitations. Specifically, contemporary DNN optimizers depend on mini-batch SGD to facilitate highly parallel training. In the Von Neumann computation architecture, the most efficient parallel mechanism is data parallelism [3], where DNN weights are replicated to different accelerators to support gradient descent with large batches [4, 5]. By contrast, in the AIMC architecture, trainable weights are stored in the analog tiles as the conductances of resistive elements; thus, copying weights is expensive and prone to error. Therefore, data parallelism, the enabling factor of large-scale AI training, is unrealistic for AIMC accelerators.
Due to the physical limitations on the AIMC accelerator, we instead resort to pipeline parallelism [6]. Pipeline parallelism partitions a large model into a series of stages distributed across multiple accelerators, with each stage corresponding to the computation of a few layers. After an accelerator processes one datum (or batch), it passes the intermediate outputs to the next accelerator and begins processing the next datum. Pipeline parallelism overlaps computation across different accelerators, increasing system throughput.
Synchronous and asynchronous parallelism. As the requirement of training large models emerges, a plethora of pipeline parallel libraries have been developed, including Gpipe [7, 8], Deepspeed [9], FSDP [10], Megetron-LM [11], Colossal-AI [12], just to name a few. In their implementation, a gradient mini-batch (e.g., with batch size 128) is split into micro-batches (e.g., 8 micro-batches with batch size 16) and filled into the pipeline. The model weights are updated after all the micro-batches are processed. Featuring the synchronization operation at the end of each mini-batch, this pipeline paradigm is referred to as synchronous pipeline. However, during the fill and drain phases, the pipeline is not fully occupied, which limits the device utilization. To overcome the issue, asynchronous pipeline is proposed, which continuously fills the micro-batch into the pipeline and updates the weights once the gradient is available. Asynchronous pipeline increases the device utilization, while the weights for gradient computing can be stale, leading to a biased training dynamic.
To fully realize the promise of analog training, this paper focuses on the theoretical understanding of the viability of stochastic gradient descent on AIMC hardware with asynchronous pipeline (Analog-SGD-AP). Besides the stale signal issue in the asynchronous pipeline, the update dynamics on AIMC accelerators also suffer from bias since the weight update involves adjusting conductance, which is difficult to control in theory. Both stale signals and imperfect update dynamics lead to the following natural question:
How do their combined effects provably impact training on the AIMC accelerator?
Despite two sources of bias, this paper theoretically demonstrates that Analog-SGD-AP converges at almost the same rate as SGD on a single AIMC accelerator (Analog-SGD-WOP) and SGD on AIMC hardware with synchronous pipeline (Analog-SGD-SP), providing a theoretical foundation for parallel AIMC training.
I-A Main results
This paper aims to understand the impact of bias and investigate the computational complexities and acceleration effect of Analog-SGD-AP. We highlight the unique technical challenges compared with the analysis of the asynchronous pipeline on digital accelerators as follows.
  1. T1)
    Unlike the digital training with precise weight update, the update dynamics on AIMC accelerators include an asymmetric bias. To be specific, consider an analog weight $W_{k}$ on a resistive crossbar array and a desired update $\Delta W_{k}$. The updated weight becomes
    $\textsf{Analog Update}\quad W_{k+1}=W_{k}+\Delta W_{k}-\frac{1}{\tau}|\Delta W_{k}|\odot W_{k}^{(m)}$ (1)
    where $\tau$ is a hardware-specific constant, $|\,\cdot\,|$ is element-wise absolute value, and $\odot$ is the element-wise production. Unlike the accurate digital training, the last bias term in (1) is unique to analog accelerators. This bias can not be eliminated within a reasonable cost and leads to challenges to the analysis. In section II-B, we provide more background.
  2. T2)
    Even though there are already theoretical works on the design of the asynchronous pipeline, most of them model it as a delayed gradient method [13]. However, in the asynchronous pipeline, the forward pass is performed using stale weights, while the backward pass uses the latest weights. Consequently, their inconsistent multiplication used for weight updates is not a valid gradient, and this error accumulates during training. The stale signals in DNN render the update increment used in Analog-SGD-AP no longer a valid gradient, making its convergence behavior unclear.
  3. T3)
    Multi-layer DNN models accumulate error through layers, leading to challenges in the analysis. Without considering a specific multi-layer architecture, prior work [14] treats all weights as a whole, thereby failing to capture the multi-layer DNN structure.
Pipeline Type Sample Complexity Comp. Density Wall-clock Complexity
Analog-SGD-WOP ${\mathcal{O}}\left(\sigma^{2}\varepsilon^{-2}\right)$ $\frac{1}{M}$ ${\mathcal{O}}\left(M\sigma^{2}\varepsilon^{-2}\right)$
Analog-SGD-SP [Thm. 1] ${\mathcal{O}}\left(\sigma^{2}\varepsilon^{-2}\right)$ $\frac{B}{M+B-1}$ ${\mathcal{O}}\left((M+B)B^{-1}\sigma^{2}\varepsilon^{-2}\right)$
Analog-SGD-AP [Thm. 2] ${\mathcal{O}}\left(\sigma^{2}\varepsilon^{-2}+\varepsilon^{-1}\right)$ 11 ${\mathcal{O}}\left(\sigma^{2}\varepsilon^{-2}+\varepsilon^{-1}\right)$
TABLE I: Comparison of different pipeline strategies. Sample complexity is the number of samples needed to achieve a training loss $\leq\varepsilon$. Computing density is the average portion of active accelerators. In the table, $M$ and $B$ are the number of stages and micro-batches, respectively. Wall-clock complexity measures the running time for achieving training loss $\leq\varepsilon$, which in math is sample complexity divided by computation density.
Overcoming the challenges, this paper answers the question above and provides both a convergence guarantee and computational complexities. The key contributions are as follows.
@@ 1. C1)
    We consider the training of a multi-layer DNN, which consists of linear layers and non-linear activation functions. We formulate the training dynamics of Analog-SGD-AP, which reflects the stale weights in the asynchronous pipeline and the bias from the AIMC accelerators’ imperfection. We investigate the smoothness of the training objective as well as the robustness of the network against stale weights.
@@ 2. C2)
[…]
@@ 3. C3)
    We verify the speedup of Analog-SGD-AP against Analog-SGD-WOP and Analog-SGD-SP in AIMC accelerator simulator, AIHWKIT. Numerical simulation results show that despite a slightly larger sample complexity, Analog-SGD-AP achieves the same accuracy with fewer clock cycles due to its high computation density, which is consistent with the theoretical guarantee. Furthermore, empirical results show that Analog-SGD-AP achieves linear speedup at least within the range of 1-8 accelerators.
I-B Prior art.
To put our work in context, we review prior art that we group into the following two categories.
Gradient-based training on AIMC accelerators. Analog training has shown promising early successes in certain tasks [15, 16, 17], and a series of analog-friendly algorithms has been proposed to enable the gradient-based on-chip training on the AIMC hardware. However, the asymmetric update and noisy gradient lead to large asymptotic errors in training [18]. Numerous algorithms have been proposed to address these issues. TT-v1 [19] heuristically introduces an auxiliary analog array as a gradient approximation to eliminate the asymptotic error. To demonstrate its efficiency, [20] establishes a model to characterize the dynamics of the analog training and provides a theoretical convergence guarantee. Building on that, [21] investigates the impact of response functions on the training and demonstrates how its impact can be alleviated by bilevel optimization methods. To filter out the high-frequency noise introduced by inter-device transfer, TT-v2 [22] introduces a digital array. Based on TT-v1/v2, Chopped-TTv2 (c-TTv2) and Analog Gradient Accumulation with Dynamic reference (AGAD) [23] are proposed to mitigate the impact of inaccurate symmetric point correction [24]. To overcome the limit update granularity issue, [25] proposed a multi-tile training paradigm. However, the mentioned works focus on training single AIMC accelerators. Instead, this work aims to scale up the system by parallelizing model computation.
Asynchronous pipeline. Existing works proposed methods to alleviate the stale weight issues in the asynchronous pipeline. For instance, [26] proposed weight-prediction approaches to compensate for the delay, PipeDream [27] combined the synchronous and asynchronous pipeline parallelisms to achieve greater speedup; and [28] and [29] propose only overlapping gradient computation and communication to ensure the delay is one. These algorithms are not for analog pipeline training. Researchers also attempted to design analog pipeline circuits to accelerate the training [30, 31], but they focus on the synchronous pipeline. On the theory side, there are few studies on pipeline training for digital accelerators such as [14, 32], but their algorithms either require weight stashing, which is expensive for analog accelerators, or slow down digital SGD through higher-order terms. As [27] points out, delay affects both forward and backward passes in the asynchronous pipeline DNN training, rendering the increment no longer a valid gradient and making the analysis challenging.
II Analog Training with Model Parallelism
In this section, we first introduce the multi-layer DNN training model, and then introduce the analog stochastic gradient descent algorithm with the asynchronous pipeline.
II-A Training problem and back-propagation
Consider a standard model training problem on a data distribution $\mathcal{P}$} given by
$\min_{W\in{\mathbb{R}}^{D}}\Big\{f(W):={\mathbb{E}}_{\xi\sim{\mathcal{P}}}[f(W;\xi)]\Big\}$ (2)
@@ II-A Training problem and back-propagation
where $W$ is the trainable model weights on analog crossbar tiles, which is vectorized into a $D$-dimension vector.
@@
DNN model. Consider the multi-layer DNN, which consists only of linear layers and activation functions. Let $M]:=\{1,2,\cdots,M\}$ denote an index set. The trainable weights are separated into $M$ different blocks, i.e., $W=[W^{(1)},W^{(2)},\cdots,W^{(M)}]$, where $W^{(m)},\forall m\in[M]$ is a matrix. The training of DNN typically consists of three phases: forward, backward, and update. Suppose $x$ and $y$ are the feature and label of data $\xi=(x,y)\in{\mathcal{X}}\times{\mathcal{Y}}$, where $\mathcal{X}$} and $\mathcal{Y}$} are feature and label spaces, respectively. The forward pass of the model for layer $m\in[M]$ is expressed by
Forward $x^{(1)}=x,~~~~z^{(m)}=W^{(m)}x^{(m)},~~~~x^{(m+1)}=g^{(m)}(z^{(m)})$ (3)
@@
where $g^{(m)}(\,\cdot\,)$ is the activation function of layer $m$, and $x^{(m)}$ is the input feature. At the last layer, the ouput $x^{(M)}$ and label $y$ are fed together into the loss function $\ell(\,\cdot\,,\,\cdot\,):{\mathcal{Y}}\times{\mathcal{Y}}\to{\mathbb{R}}$ and yields $f(W;\xi)=\ell(x^{(M+1)},y)$. For the backward pass, define the backward signal $\delta^{(m)}:=\frac{\partial f(W;\xi)}{\partial z^{(m)}}$, wherein the error of the last layer is first computed by $\delta^{(M)}=\nabla_{x^{(M+1)}}\ell(x^{(M+1)},y)\nabla g^{(M)}(z^{(M)})$. Subsequently, each layer $m\in[M-1]$ computes the next recursion by the chain rule
@@
In the forward pass, the activation $z^{(m)}$ is stashed until the backward signal $\delta^{(m+1)}$ returns. Typically, the computation $\nabla g^{(m)}(z)$ is light-weight, and hence the major overhead of the backward pass lies at the MVM among $W^{(m)}$ and $\delta^{(m)}$. With the forward and backward signals, the gradient with respect to the weight $W^{(m)}$ of layer $m$ is computed by
@@
where $\otimes$ is the outer product between two vectors.
Rank-update Unlike digital accelerators, AIMC accelerators use rank-update to change the matrix, which leverages the outer product structure as shown in (5). To do so, independent random voltage pulses with probability proportional to the elements of $\delta^{(m)}$ and $x^{(m)}$ are applied to columns and rows, respectively. Consequently, the $(i,j)$-th element $W]_{i,j}$ receives a pulse coincident with probability $\propto[\delta^{(m)}]_{i}[x^{(m)}]_{j}$. Without computing the whole matrix $\delta^{(m)}\otimes x^{(m)}$ explicitly, rank-update has only $O(BN)$ complexity while the update on digital accelerators has $O(BN^{2})$ complexity to update an $N\times N$ matrix. Consequently, AIMC accelerators perform the update operation in a fast and energy-efficient way [15].
@@
Figure 2: Illustration of pipelines with 4 accelerators ($M=4$). Each mini-batch is split into $B=4$ micro-batches. Each color represents one micro-batch, and each row from bottom to top represents stages 0 to 4. Each column corresponds to a single clock cycle in which one micro-batch is processed. The white square indicates the corresponding accelerator is idle. (Top) Analog-SGD-WOP. All weights are updated after each mini-batch is computed, which is not presented in the figure. (Middle) Analog-SGD-SP with micro-batch count $B=5$. The update occurs after all 5 micro-batches are processed. (Bottom) Analog-SGD-AP. The update occurs once the gradient of a single micro-batch is achieved. The grey square indicates that the corresponding accelerator is processing data that is not fully reflected in the figure.
II-B Non-ideal update on AIMC accelerator
Despite the ultra-efficient weight update on the AIMC accelerator, as noted by [19, 20, 21], the update on AIMC accelerators exhibits asymmetric bias. To be specific, for any desired update $\Delta W_{k}^{(m)}$, the update on $W_{k}$ is scaled by response functions $q_{+}(\,\cdot\,)$ and $q_{-}(\,\cdot\,)$. Considering the weight $W_{k}$ at iteration $k$ and the expected update $\Delta W\in{\mathbb{R}}^{D}$, we express the asymmetric update as $W_{k+1}=U(W_{k},\Delta W)$ with the update function $U:{\mathbb{R}}\times{\mathbb{R}}\to{\mathbb{R}}$ defined by
$U(w,\Delta w):=\begin{cases}w+\Delta w\cdot q_{+}(w),~~~\Delta w\geq 0,\\ w+\Delta w\cdot q_{-}(w),~~~\Delta w<0.\end{cases}$ (6)
Defining the symmetric and asymmetric components as $F(w):=\frac{1}{2}(q_{-}(w)+q_{+}(w))$ and $G(w):=\frac{1}{2}(q_{-}(w)-q_{+}(w))$, we write the update in (6) as $U(w,\Delta w)=w+\Delta w\cdot F(w)-|\Delta w|\cdot G(w)$. This paper consider a special case that $q_{+}(w)=1-{w}/{\tau},q_{-}(w)=1+{w}/{\tau}$ as suggested by [20], we reach the dynamics (1).
Figure 3: Illustration of the dynamics of the asynchronous pipeline. The $W_{k}^{(m)}$ in the circle implies the update happens in this clock cycle, and the symbols in the squares indicate the input of each accelerator.
II-C Synchronous and asynchronous pipelines